# BlockFusion: Expandable 3D Scene Generation using Latent Tri-plane Extrapolation

### v3

@@ 4.4. Ablation study
Shape reconstruction quality: latent tri-plane vs raw-tri-plane. Figure 10 and Table 3 display the qualitative and quantitative room block reconstruction results using different representations. The raw tri-plane can accurately represent the ground truth (GT) mesh without any issues. Compared to the raw tri-plane, the latent tri-planes with 2 channels manage to reduce $99.6\%$ of the data bits while still maintaining decent shape representation power. Using a similar data compression rate, the 4096-dimensional latent vector cannot produce any reasonable shape. Considerably, the raw tri-plane is a redundant 3D representation. However, when we attempted to use fewer feature channels during the raw tri-plane fitting, we found that the shape reconstruction quality was significantly degraded. This observation demonstrates the necessity of using an auto-encoder for tri-plane compression.
Table 3. Quantitative reconstruction results for indoor blocks. CPR: compression rate w.r.t the raw tri-plane. The units are CD with scale $10^{-3}$ , $E_{NRM}$ in degrees, and $E_{SDF}$ in centimeters.
@@
Shape generation quality: latent tri-plane vs raw-tri-plane. Figure 4 and Table 2 display the qualitative and quantitative unconditional room block generation results. The latent tri-plane shows significantly better results, with a $27.84\%$ and $25.83\%$ increase in Coverage scores under the CD and EMD metrics respectively. Qualitatively, raw tri-plane diffusion can not produce any reasonable results. In conclusion, compared to the raw tri-plane, the latent tri-plane retains decent shape representation capacity while serving as a superior proxy for shape generation.
@@
How does resampling affect the extrapolation? Fig. 12 shows the shape synchronization results of layout-conditioned tri-plane extrapolation. We tested different resampling times with $R=\{1,2,3,7\}$. The chamfer distance drops steadily with more resampling steps and stabilizes after 3 resamplings, where the variance in the Chamer Distance also converges. This suggests that augmenting the number of resampling times can improve the quality of synchronization results. For clarity, $R=0$ means that we do not perform synchronizations, i,e. the two blocks are generated independently while adhering to the shared layout conditions. Note that in this case, the Chamfer Distance is extremely high, indicating that using layout conditioning alone does not ensure consistent geometry between blocks.

### v4

@@ 4.4. Ablation study
Shape reconstruction quality: latent tri-plane vs raw-tri-plane. Figure 10 and Table 4 display the qualitative and quantitative room block reconstruction results using different representations. The raw tri-plane can accurately represent the ground truth (GT) mesh without any issues. Compared to the raw tri-plane, the latent tri-planes with 2 channels manage to reduce $99.6\%$ of the data bits while still maintaining decent shape representation power. Using a similar data compression rate, the 4096-dimensional latent vector cannot produce any reasonable shape. Considerably, the raw tri-plane is a redundant 3D representation. However, when we attempted to use fewer feature channels and resolution during the raw tri-plane fitting, we observed a considerable decline in the quality of shape reconstruction, as depicted in the first two rows of Table 4. This observation demonstrates the necessity of using an auto-encoder for tri-plane compression.
Table 4. Quantitative reconstruction results for indoor blocks. CPR: compression rate w.r.t the raw tri-plane at resolution $3\times 128^{2}\times 32$ The units are CD with scale $10^{-3}$ , $E_{NRM}$ in degrees, and $E_{SDF}$ in centimeters.
@@
Shape generation quality: latent tri-plane vs raw-tri-plane. Figure 4 and Table 3 display the qualitative and quantitative unconditional room block generation results. The latent tri-plane diffusion ($z_{0}$) shows significantly better results, with a $27.84\%$ and $25.83\%$ increase in Coverage scores under the CD and EMD metrics respectively. Qualitatively, raw tri-plane diffusion can not produce any reasonable results. In conclusion, compared to the raw tri-plane, the latent tri-plane retains decent shape representation capacity while serving as a superior proxy for shape generation.
@@
How does resampling affect the extrapolation? Fig. 12 shows the shape synchronization results of layout-conditioned tri-plane extrapolation. We tested different resampling times with $R=\{1,2,3,7\}$. The Chamfer Distance drops steadily with more resampling steps and stabilizes after 3 resamplings, where the variance in the Chamfer Distance also converges. This suggests that augmenting the number of resampling times can improve the quality of synchronization results. For clarity, $R=0$ means that we do not perform synchronizations, i,e. the two blocks are generated independently while adhering to the shared layout conditions. Note that in this case, the Chamfer Distance is extremely high, indicating that using layout conditioning alone does not ensure consistent geometry between blocks.
@@
Predicting $z_{0}$ vs Predicting noise. In Section 3.4, we adapt the strategy that predicts $z_{0}$ with $\Psi$ during the reverse diffusion process. This is in contrast to the vanilla DDPM (Ho et al., 2020) which predict the noise. We conduct a quantitative ablation study to compare the two strategies. As shown in Table 3, predicting $z_{0}$ and predicting noise as targets achieves comparable results in terms of unconditional generation.
@@ 5. Conclusion and Discussion
Appendix A Qualitative room generation
To reduce variables in the experiment, we validated the coloring of generated geometry by Meshy for Text2room (Höllein et al., 2023). Additional qualitative and quantitative results are presented in Fig. 15 and Table 5. A total of 9 people are included in this user study.
Figure 15. Supplementary room generation results. Text2Room (Höllein et al., 2023) generates distorted shapes. While using Meshy to generate texture, it cannot correspond well to the input prompts.
Table 5. Quantitative indoor scene generation results of textured shape.
[…]
Text2Room + Meshy 1.22 1.22
Ours + Meshy 4.56 4.67
Appendix B VAE Structure
Fig. 16 and 17 details our VAE for compressing raw tri-plane to latent tri-plane space as described in Section 3.3. Parameters of ResNet (He et al., 2016) blocks are given as (#in channels, #out channels). Parameters of Transformer (Dosovitskiy et al., 2020) layers are given as (# channels) All shapes are denoted as (#batch,#channel,Height,Width). $C$ refers to the number of channels compressed into the latent tri-plane space. Feature Pyramid Networks (Lin et al., 2017) are hired to aggregate multi-scale features. The tri-planes are unfolded into three independent planes to run convolutions separately. Transformer layers are leveraged to achieve cross-plane dependence.
Figure 16. VAE structure. VAE enocder structure.
Figure 17. VAE structure. VAE decoder structure.